OpenStudy (anonymous):

Completely factor and show all steps: y^3-(x+1)^3

OpenStudy (anonymous):

a^3-b^3=(a-b)(a^2+ab+b^2)

OpenStudy (anonymous):

@ghass1978 It isn't an equation. You just need to factor it, which I don't know how to do.

OpenStudy (anonymous):

thats what i did...y is a and (x+1) is b

OpenStudy (anonymous):

How come you added an equals sign?

OpenStudy (anonymous):

if you factorize an expression you get an equal expression y^3-(x+1)^3=(y-x-1)(y^2+y(x+1)+(x+1)^2)